# Imprints of the Local Bubble and Dust Complexity on Polarized Dust Emission

### v1

@@ 4.5 Results
Figure 8: The dust distribution in units of differential extinction (mag/pc, left y-axis) and equivalent Hydrogen number density (cm-3, right y-axis) along a pair of matched sightlines through the first posterior sample of the Edenhofer et al. (2023) map. These sightlines have the same $N_{\rm H}^{\rm Edenhofer}$, but the top one has a higher complexity than the bottom one. The Galactic coordinates, dust complexity, and polarization fraction of each of the sightlines are denoted on their subpanels.
@@
We perform the analysis described in the previous subsections on the Edenhofer et al. (2023) maps which extend radially to 1.25 kpc and compare the polarization fractions of sightlines with $\mathcal{N}_{c}\leq 1.1$ and those with $\mathcal{N}_{c}\geq 1.5$. An example of a pair of matched sightlines with the same $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ is shown in Figure 8. Even though the polarization fraction of the higher-complexity sightline is lower than that of the lower-complexity sightline in this example, not all sightlines follow this trend. We are only looking for a statistically significant average effect.
@@
For this test, $\Delta p_{353}^{s}$, the mean over the paired sightlines for each posterior sample $s$, is plotted in orange at the top of Figure 9. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 samples are $-1.47~{}\times~{}10^{-2}$ and $0.22~{}\times~{}10^{-2}$, respectively. All 12 posterior samples pass the permutation test described in Section 4.4 with a p-value $<0.001$. A random $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample is shown in blue at the top of Figure 9 as well. Therefore, we find that higher dust complexity at equivalent column densities is associated with depolarization at the 1.5% level. This is at the level of 6.8% of the maximum polarization fraction measured by Planck Collaboration et al. (2020a) at 353 GHz and 80.
@@
To determine whether this result is uniquely enabled by the Edenhofer et al. (2023) dust maps since they extend radially to 1.25 kpc, we repeat the analysis using the 3D Leike et al. (2020) dust maps, which extend to 370 pc in the positive and negative Galactic-X and Y coordinates and 270 pc in the positive and negative Galactic-Z coordinate. For consistency between different sightlines in those maps, we truncate all sightlines at 270 pc. Also, we start each sightline at 70 pc since Leike et al. (2020) find that the reconstructed dust density closer than 70 pc resembles a smeared-out version of the farther dust, an artifact related to systematic data biases. Leike et al. (2020) also provide 12 posterior samples for their maps which we use for this analysis.
@@
The results for repeating the analysis using the Leike et al. (2020) maps instead are shown at the bottom of Figure 9. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 posterior samples are $2.75~{}\times~{}10^{-3}$ and $3.06\times~{}10^{-3}$, respectively. None of the 12 samples pass the permutation test, i.e., they are indistinguishable from the distributions of $\Delta p_{353}^{s,\,{\rm null}}$. We also show a random $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample in blue in the same subplot of Figure 9.
@@
To determine whether the null result is attributed to using a different dataset or to the lower extent in radial distance, we run the analysis on the Edenhofer et al. (2023) maps up to 270 pc, the same distance used for the Leike et al. (2020) maps. We plot the results in the middle of Figure 9. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 posterior samples are $1.05~{}\times~{}10^{-3}$ and $1.36~{}\times~{}10^{-3}$, respectively. Again, none of the 12 samples pass the permutation test, i.e., they are indistinguishable from the distributions of $\Delta p_{353}^{s,\,{\rm null}}$. The consistency of this result with null as well indicates that the null result we found using the Leike et al. (2020) maps is due to only considering distances up to 270 pc, not the choice of 3D dust dataset. This illustrates the role that dust components farther than 270 pc play in affecting polarization measurements.
@@
Figure 9: The mean difference in $p_{353}$ over the paired sightlines for each of the 12 posterior samples, where for each pair, $p_{353}$ for the lower-complexity sightline is subtracted from $p_{353}$ for the higher-complexity sightline. $\mathcal{N}_{c}$ is calculated on the Edenhofer et al. (2023) maps up to 1.25 kpc (top), the Edenhofer et al. (2023) maps up to 270 pc (middle), and the Leike et al. (2020) maps up to 270 pc (bottom). For each panel, the mean differences for the actual test are plotted in orange, and those for the null test are plotted in blue. The mean and standard deviation of the 12 means for each test are also plotted. The top panel agrees with our hypothesis: that sightlines with similar column densities will, on average, exhibit lower dust polarization fractions when their 3D dust distribution is more complex.
@@ 4.6 Validation
In this subsection, we summarize some of the tests we performed to verify our results. As described in Section 4.2, we find that our results are robust to reasonable variations in $\Delta_{n}$ and #$\#_{\rm voxels}$. The results are also consistent when using the Planck R3.01 353 GHz maps instead of the GNILC maps.
To ensure that the pairs were matched correctly, we examine the distributions of the differences in $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ between the matched pairs. We verify that those differences peak near zero and are not skewed toward the positive or negative directions. We find that to be the case for both variables and over all 12 posterior samples of Edenhofer et al. (2023). We also find our results to be robust when including $|b|$ as a confounding variable in addition to $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ in the pair matching – in other words, the result is not attributable to a dependence of the polarization fraction on Galactic latitude.
@@
We experiment with varying the $\mathcal{N}_{c}$ thresholds that define the bin edges of the low- and high-complexity sightlines. The 25th and 75th percentiles of the $\mathcal{N}_{c}$ distribution over our mask vary slightly over the 12 posterior samples of Edenhofer et al. (2023) and over variations in $\Delta_{n}$ and #$\#_{\rm voxels}$. However, they are roughly 1.4 and 2.3, respectively. Therefore, we split our sightlines based on these values into low- ($\mathcal{N}_{c}\leq 1.4$) and high-complexity ($\mathcal{N}_{c}\geq 2.3$) bins to achieve a roughly equal number of sightlines in each group. This should improve the pair-matching outcomes since there are more sightlines to match from in the smaller group. We find that our results are robust to this change.
Since neighboring sightlines are likely to have similar $\mathcal{N}_{c}$ and similar $p$ values, we test whether large regions of neighboring sightlines belonging to either the low- or high-complexity groups bias our results. We randomly sample 1,000 sightlines out of about 9,000 sightlines from each of the the low- and high-complexity groups before pair matching. We find that our results are robust to this test.
@@
i.e., we use the total column density out to 1.25 kpc rather than a sum over the dust components in the numerator. With the new definition, $\mathcal{N}_{c}$ is only sensitive to the dendrogram-identified component with the highest column density $N^{\rm max}_{\rm H}$ rather than also being sensitive to the other dendrogram-identified components. Since the dendrogram parameters define the minimum prominence for a peak to be considered an independent component, we expect $N^{\rm max}_{\rm H}$ to be the least sensitive component to those parameters. Therefore, the new definition is much less sensitive to the dendrogram parameters. We test whether our results are robust to using the new definition and confirm that they are.
@@ 5 Discussion and Conclusions
In this paper, we explore how different geometrical factors affect the polarization fraction. In Section 3, we test whether we detect an imprint of the Local Bubble geometry on the polarization fraction. Following the well-motivated assumptions that the magnetic field lines are tangential to the Local Bubble surface and that this surface is defined by the model of Pelgrims et al. (2020), we test for a correlation between the measured polarization fraction and the angle the line of sight makes with the tangential magnetic field lines. However, we do not find evidence for this in sightlines where the dust extinction is dominated by the Local Bubble. We also do not find a correlation between the polarization fraction and the inclination angle between the Local Bubble wall defined by the model of O’Neill et al. (2024) and the plane of the sky. Therefore, we conclude that at least one of the commonly made assumptions must not hold. We hypothesize that dust structure beyond the Local Bubble wall plays a substantial role in determining the polarization structure of the dust emission. Our results show that simply projecting the Planck polarization data onto the Local Bubble geometry is not a well-motivated model for the magnetic field structure of the Local Bubble.
@@
In Section 4, we test how dust complexity, i.e., how the 3D dust is distributed along the line of sight, affects the polarization fraction. We quantify the dust complexity for each sightline and group the sightlines into low- and high-complexity groups. We pair-match the sightlines across the two groups based on their column densities. For each pair, we subtract the polarization fraction of the sightline with low complexity from that of the sightline with high complexity. We find that on average, the polarization fraction of the sightlines with higher complexity is 2% lower than those with lower complexity. This is only true when considering dust out to 1.25 kpc. The result is not statistically different from null when considering dust out to 270 pc only. Note that our definition of complexity does not take into account the distance to different dust components. Future work could incorporate the effect of this distance into the analysis.
@@
We test whether the order of magnitude of this result agrees with our expectation based on geometric depolarization. Padoan et al. (2001) model the polarized thermal dust emission from protostellar cores formed through supersonic turbulent flows within molecular clouds following the formalism in Fiege & Pudritz (2000). Fiege & Pudritz (2000) develops this formalism to model the submillimeter polarization patterns for filamentary molecular clouds. Fiege & Pudritz (2000) and Padoan et al. (2001) ignore the effects of self-absorption and scattering since this model is for submillimeter wavelengths. Padoan et al. (2001) further assume that the dust grain properties are constant and the temperature is uniform. Since these assumptions are valid for our order of magnitude estimation, we follow the same formalism here.
Combining Equations 12, 13, and 14 of Padoan et al. (2001), the polarization fraction for a given sightline can be written as
[…]
where $q~{}=~{}Q/I$$u~{}=~{}U/I$$\rho$ is the volume density, $ds$ is a distance segment along the line of sight, $\alpha$ is a coefficient defined in Equation 15 of Padoan et al. (2001) that is a product of polarization reduction factors, such as the degree of dust alignment, and the dust grain polarization cross-section, and $\gamma$ is the angle between the magnetic field vector and the plane of the sky.
From Equations 8 and 9 of Padoan et al. (2001), we have
$q~{}=~{}\int\rho\cos{2\psi}\cos^{2}{\gamma}\,ds,$ (13)
$u~{}=~{}\int\rho\sin{2\psi}\cos^{2}{\gamma}\,ds,$ (14)
where $\psi$ is the angle between the projection of the magnetic field on the plane of the sky and North.
@@
For this test, we consider sightlines with 2 clouds but different $\mathcal{N}_{c}$ values. For simplicity, we assume each cloud to have a constant volume density, plane-of-sky magnetic field angle, and magnetic inclination angle along a certain sightline. Therefore, for a given sightline, we write
$q~{}=~{}N_{\rm H,a}\cos{2\psi_{1}}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\cos{2\psi_{2}}\cos^{2}{\gamma_{\rm b}},$ (15)
$u~{}=~{}N_{\rm H,a}\sin{2\psi_{1}}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\sin{2\psi_{2}}\cos^{2}{\gamma_{\rm b}},$ (16)
@@
where $N_{\rm H,a}$ and $N_{\rm H,b}$ are the column densities for each cloud, and $\gamma_{\rm a}$ and $\gamma_{\rm b}$ are the magnetic inclination angles of the two clouds. Similarly, Equation 12 can be written as
$p=\frac{0.15\sqrt{q^{2}+u^{2}}}{N_{\rm H}-\frac{0.15}{2}\left(N_{\rm H,a}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\cos^{2}{\gamma_{\rm b}}-\frac{2}{3}N_{\rm H}\right)},$ (17)
where we set $\alpha=0.15$ as in Padoan et al. (2001), which corresponds to a $p_{\rm max}=0.15$.
@@
and divide the numerator and denominator of Equation 17 by $N_{\rm H,a}$ to be able to write it in terms of $\mathcal{N}_{c}$, where  $N_{\rm H}/N_{\rm H,a}~{}=~{}\mathcal{N}_{c}$ and $N_{\rm H,b}/N_{\rm H,a}~{}=~{}1~{}-~{}\mathcal{N}_{c}$.
@@
We randomly sample an orientation for $\psi_{1}$$\psi_{2}$$\gamma_{\rm a}$, and $\gamma_{\rm b}$ and a value between 1 and 1.1 for $\mathcal{N}_{c}$ to calculate an instance of $p(\mathcal{N}_{c}\leq 1.1)$. We also sample different orientations and a value between 1.5 and 2 for $\mathcal{N}_{c}$ to calculate an instance of $p(\mathcal{N}_{c}\geq 1.5)$. We then subtract $p(\mathcal{N}_{c}\leq 1.1)$ from $p(\mathcal{N}_{c}\geq 1.5)$ as in Section 4. We repeat this 10,000 times and average the results. We get
$\langle p(\mathcal{N}_{c}~{}\geq~{}1.5)~{}-~{}p(\mathcal{N}_{c}~{}\leq~{}1.1)\rangle~{}=~{}-0.015.$ (19)
@@
This agrees with the result we measure in Section 4. Thus our empirical result is consistent with our theoretical estimate for the dust depolarization attributable to the line-of-sight dust complexity.
@@
The line-of-sight magnetic field component is not directly measurable from the dust polarization. Variations in the orientation of magnetic fields along the line of sight induce differences in the polarization angles of different dust components along the same sightline (Lee & Draine, 1985; Tassis & Pavlidou, 2015; King et al., 2018). When the emission of those components has different spectral energy distributions (SEDs), a frequency-dependent variation of the observed dust polarization angle along that sightline emerges, a phenomenon known as line-of-sight frequency decorrelation. This decorrelation complicates the translation of polarized dust emission maps from one frequency to another. Current analysis within the BICEP/Keck field does not demonstrate evidence of this phenomenon (BICEP/Keck Collaboration et al., 2021, 2023). However, a statistically significant detection of line-of-sight frequency decorrelation has been identified in larger sky areas across sightlines intersecting multiple dust clouds with varying magnetic field orientations (Pelgrims et al., 2021). Since polarized dust emission is the major foreground for CMB polarization measurements at high frequencies, it is important to characterize how the spatial complexity of the magnetic field in the dust might affect the frequency dependence of the foreground signal. In this paper, we have presented evidence that the spatial complexity of the dust affects the dust polarization signal.
@@
The analysis in this paper highlights the importance of 3D dust mapping out to large distances. The dust distribution affects the dust polarization fraction, which has implications for the 3D magnetic field distribution. Since we expect higher complexity sightlines to have a lower polarization fraction on average, we would know that the magnetic field is more uniform along a sightline if it has both a highly complex dust distribution and a large polarization fraction. This analysis was performed on the sightlines shown in Figure 5. Improvements to 3D dust modeling will allow us to look for this effect at the very high Galactic latitudes excluded here.

### v2

@@ 4.5 Results
Figure 9: The dust distribution in units of differential extinction (mag/pc, left y-axis) and equivalent Hydrogen number density (cm-3, right y-axis) along a pair of matched sightlines through the first posterior sample of the Edenhofer et al. (2023) map. These sightlines have the same $N_{\rm H}^{\rm Edenhofer}$, but the top one has a higher complexity than the bottom one. The Galactic coordinates, dust complexity, and dust polarization fraction of each of the sightlines are denoted on their subpanels.
Figure 10: Kernel density estimate plots of the $\Delta p_{353}$ distributions over all matched pairs of sightlines of one posterior sample of the Edenhofer et al. (2023) maps. The $p_{353}$ of the lower complexity sightline is always subtracted from that of the higher complexity one in the orange distribution. This has a mean of $\Delta p_{353}=-1.81~{}\times~{}10^{-2}$, which is plotted as an orange dot in Figure 11. The blue distribution contains the same pairs as the orange one with the sign randomly flipped for each pair, i.e., the distribution of the permutation-based null test described in Section 4.4.
@@
We perform the analysis described in the previous subsections on the Edenhofer et al. (2023) maps which extend radially to 1.25 kpc and compare the dust polarization fractions of sightlines with $\mathcal{N}_{c}\leq 1.1$ and those with $\mathcal{N}_{c}\geq 1.5$. An example of a pair of matched sightlines with the same $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ is shown in Figure 9. Even though the dust polarization fraction of the higher-complexity sightline is lower than that of the lower-complexity sightline in this example, not all sightlines follow this trend. We are only looking for a statistically significant average effect. An example of the distribution of the differences in $p_{353}$ over all pairs in one of the map posterior samples is shown in Figure 10 along with the same distribution for the permutation-based null test.
@@
For this test, $\Delta p_{353}^{s}$, the mean over the paired sightlines for each posterior sample $s$, is plotted in orange at the top of Figure 11. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 samples are $-1.47~{}\times~{}10^{-2}$ and $0.22~{}\times~{}10^{-2}$, respectively. All 12 posterior samples pass both null tests described in Section 4.4 with a p-value $<0.001$. A random $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample is shown in blue at the top of Figure 11 as well. Therefore, we find that higher dust complexity at equivalent column densities is associated with depolarization at the 1.5% level. This is at the level of 6.8% of the maximum dust polarization fraction measured by Planck Collaboration et al. (2020a) at 353 GHz and 80.
@@
To determine whether this result is uniquely enabled by the Edenhofer et al. (2023) dust maps since they extend radially to 1.25 kpc, we repeat the analysis using the 3D Leike et al. (2020) dust maps, which extend to 370 pc in the positive and negative Galactic-X and Y coordinates and 270 pc in the positive and negative Galactic-Z coordinate. For consistency between different sightlines in those maps, we truncate all sightlines at 270 pc. Also, we start each sightline at 70 pc since Leike et al. (2020) find that the reconstructed dust density closer than 70 pc resembles a smeared-out version of the farther dust, an artifact related to systematic data biases. Leike et al. (2020) also provide 12 posterior samples for their maps which we use for this analysis. Because we sample these data in increments of 2 pc as opposed to 7 pc as in the case of the Edenhofer et al. (2023) data, we set the dendrogram parameter #$\#_{\rm voxels}~{}=~{}10$, which corresponds to 20 pc, compared to #$\#_{\rm voxels}~{}=~{}3$, which corresponds to 21 pc in the case of the Edenhofer et al. (2023) data. We keep the dendrogram parameter $\Delta_{n}$ the same for both maps. This results in about 3,500 sightlines with $\mathcal{N}_{c}\leq 1.1$ and about 8,500 with $\mathcal{N}_{c}\geq 1.5$.
@@
The results for repeating the analysis using the Leike et al. (2020) maps instead are shown at the bottom of Figure 11. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 posterior samples are $2.75~{}\times~{}10^{-3}$ and $3.06\times~{}10^{-3}$, respectively. None of the 12 samples pass either of the null tests described in Section 4.4, i.e., they are indistinguishable from the distributions of $\Delta p_{353}^{s,\,{\rm null}}$. We also show a random $\Delta p_{353}^{s,\,{\rm null}}$ for each posterior sample in blue in the same subplot of Figure 11.
@@
To determine whether the null result is attributed to using a different dataset or to the lower extent in radial distance, we run the analysis on the Edenhofer et al. (2023) maps up to 270 pc, the same distance used for the Leike et al. (2020) maps. The number of sightlines with $\mathcal{N}_{c}\leq 1.1$ and $\mathcal{N}_{c}\geq 1.5$ after this distance cut are about 10,000 and 3,500, respectively. The difference in the $\mathcal{N}_{c}$ distributions between this data and the Leike et al. (2020) maps is due to several differences in the maps and the post-processing we perform on them, including having to use slightly different dendrogram parameters and not counting components whose peak is closer than the distance where the differential extinction integrated radially outwards reaches 50 mmag in the Edenhofer et al. (2023) maps among other differences. We plot the results in the middle of Figure 11. The mean and standard deviation of $\Delta p_{353}^{s}$ over the 12 posterior samples are $1.05~{}\times~{}10^{-3}$ and $1.36~{}\times~{}10^{-3}$, respectively. Again, none of the 12 samples pass either of the null tests described in Section 4.4, i.e., they are indistinguishable from the distributions of $\Delta p_{353}^{s,\,{\rm null}}$. The consistency of this result with null as well indicates that the null result we found using the Leike et al. (2020) maps is due to only considering distances up to 270 pc, not the choice of 3D dust dataset. This illustrates the role that dust components farther than 270 pc play in affecting polarization measurements.
@@
Figure 11: The mean difference in $p_{353}$ over the paired sightlines for each of the 12 posterior samples, where for each pair, $p_{353}$ for the lower-complexity sightline is subtracted from $p_{353}$ for the higher-complexity sightline. $\mathcal{N}_{c}$ is calculated on the Edenhofer et al. (2023) maps up to 1.25 kpc (top), the Edenhofer et al. (2023) maps up to 270 pc (middle), and the Leike et al. (2020) maps up to 270 pc (bottom). For each panel, the mean differences for the actual test are plotted in orange, and samples from the permutation-based null tests are plotted in blue. The mean and standard deviation of the 12 means for each test are also plotted. The top panel agrees with our hypothesis: that sightlines with similar column densities will, on average, exhibit lower dust polarization fractions when their 3D dust distribution is more complex.
@@ 4.6 Validation
In this subsection, we summarize some of the tests we performed to verify our results. As described in Section 4.2, we find that our results are robust to reasonable variations in $\Delta_{n}$ and #$\#_{\rm voxels}$. These are summarized in Table 1 in Appendix A. The results are also consistent when using the Planck R3.01 353 GHz maps instead of the GNILC maps.
To ensure that the pairs were matched correctly, we examine the distributions of the differences in $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ between the matched pairs. We verify that those differences peak near zero and are not skewed toward the positive or negative values. We find that to be the case for both variables and over all 12 posterior samples of Edenhofer et al. (2023). We also find our results to be robust when including $|b|$ as an additional confounding variable to $N_{\rm H}^{\rm Edenhofer}$ and $N_{\rm H}^{\rm Planck}$ in the pair matching – in other words, the result is not attributable to a dependence of the dust polarization fraction on Galactic latitude. We show these results in Table 2 in Appendix A.
@@
We experiment with varying the $\mathcal{N}_{c}$ thresholds that define the bin edges of the low- and high-complexity sightlines. The 25th and 75th percentiles of the $\mathcal{N}_{c}$ distribution over our mask vary slightly over the 12 posterior samples of Edenhofer et al. (2023) and over variations in $\Delta_{n}$ and #$\#_{\rm voxels}$. However, they are roughly 1.4 and 2.3, respectively. Therefore, we split our sightlines based on these values into low- ($\mathcal{N}_{c}\leq 1.4$) and high-complexity ($\mathcal{N}_{c}\geq 2.3$) bins to achieve a roughly equal number of sightlines in each group. This should improve the pair-matching outcomes since there are more sightlines to match from in the smaller group. We find that our results are robust to this change as shown in Table 1 in Appendix A.
Since neighboring sightlines are likely to have similar $\mathcal{N}_{c}$ and similar $p_{353}$ values, we test whether large regions of neighboring sightlines belonging to either the low- or high-complexity groups bias our results. We randomly sample 1,000 sightlines out of about 5,000 sightlines from each of the low- and high-complexity groups before pair matching, where we use $\mathcal{N}_{c}\leq 1.4$ and $\mathcal{N}_{c}\geq 2.3$ for these groups, respectively, in this case. We find that our results are robust to this test as shown in Table 3 in Appendix A.
@@
i.e., we use the total column density out to 1.25 kpc rather than a sum over the dust components in the numerator. With the new definition, $\mathcal{N}_{c}$ is only sensitive to the dendrogram-identified component with the highest column density $N^{\rm max}_{\rm H}$ rather than also being sensitive to the other dendrogram-identified components. Since the dendrogram parameters define the minimum prominence for a peak to be considered an independent component, we expect $N^{\rm max}_{\rm H}$ to be the least sensitive component to those parameters. Therefore, the new definition is much less sensitive to the dendrogram parameters. The 25th and 75th percentiles of this modified version of $\mathcal{N}_{c}$ are about 0.05 and 0.09 for the 12 posterior samples, and we use these as the upper and lower thresholds for the low- and high-complexity bins, respectively. Over the 12 posterior samples, we find a mean and standard deviation of $\Delta p_{353}~{}=~{}0.96~{}\times~{}10^{-2}~{}\pm~{}0.34~{}\times~{}10^{-2}$. This passes the null test with a p-value${}<~{}0.001$.
@@ 5 Discussion and Conclusions
In this paper, we explore how different geometrical factors affect the fractional polarization of the dust emission. In Section 3, we test whether we detect an imprint of the Local Bubble geometry on the dust polarization fraction. Following the well-motivated assumptions that the magnetic field lines are tangential to the Local Bubble surface and that this surface is defined by the model of Pelgrims et al. (2020), we test for a correlation between the measured dust polarization fraction and the angle the line of sight makes with the tangential magnetic field lines. However, we do not find evidence for this in sightlines where the dust extinction is dominated by the Local Bubble. We also do not find a correlation between the dust polarization fraction and the inclination angle between the Local Bubble wall defined by the model of O’Neill et al. (2024) and the plane of the sky. Therefore, we conclude that at least one of the commonly made assumptions must not hold. We hypothesize that dust structure beyond the Local Bubble wall plays a substantial role in determining the polarization structure of the dust emission. Our results show that simply projecting the Planck polarization data onto the Local Bubble geometry is not a well-motivated model for the magnetic field structure of the Local Bubble.
@@
In Section 4, we test how dust complexity, i.e., how the 3D dust is distributed along the line of sight, affects the dust polarization fraction. We quantify the dust complexity for each sightline and group the sightlines into low- and high-complexity groups. We pair-match the sightlines across the two groups based on their column densities. For each pair, we subtract the dust polarization fraction of the sightline with low complexity from that of the sightline with high complexity. We find that on average, the dust polarization fraction of the sightlines with higher complexity is 2% lower than those with lower complexity. This is only true when considering dust out to 1.25 kpc. The result is not statistically different from null when considering dust out to 270 pc only. Note that our definition of complexity does not take into account the distance to different dust components. Future work could incorporate the effect of this distance into the analysis.
@@
We test whether the order of magnitude of this result agrees with our expectation based on geometric depolarization. Padoan et al. (2001) model the polarized thermal dust emission from protostellar cores formed through supersonic turbulent flows within molecular clouds following the formalism in Fiege & Pudritz (2000). Fiege & Pudritz (2000) develops this formalism to model the submillimeter polarization patterns for filamentary molecular clouds. Fiege & Pudritz (2000) and Padoan et al. (2001) ignore the effects of self-absorption and scattering since this model is for submillimeter wavelengths at which the diffuse interstellar medium is optically thin. Padoan et al. (2001) further assume that the dust grain properties are constant and the temperature is uniform. Since these assumptions are valid for our order of magnitude estimation, we follow the same formalism here.
@@
For this test, we consider sightlines with 2 clouds but different $\mathcal{N}_{c}$ values. For simplicity, we assume each cloud to have a constant volume density, plane-of-sky magnetic field angle, and magnetic inclination angle along a certain sightline. Therefore, for a given sightline, we write Equations 2,  3, and 4 as
$I~{}=\epsilon N_{\rm H}-\frac{\alpha\epsilon}{2}\left(N_{\rm H,a}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\cos^{2}{\gamma_{\rm b}}-\frac{2}{3}N_{\rm H}\right),$ (12)
$Q~{}=~{}-\alpha\epsilon(N_{\rm H,a}\cos{2\psi_{a}}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\cos{2\psi_{b}}\cos^{2}{\gamma_{\rm b}}),$ (13)
$U~{}=~{}-\alpha\epsilon(N_{\rm H,a}\sin{2\psi_{a}}\cos^{2}{\gamma_{\rm a}}+N_{\rm H,b}\sin{2\psi_{b}}\cos^{2}{\gamma_{\rm b}}),$ (14)
@@
where $N_{\rm H,a}$ and $N_{\rm H,b}$ are the column densities for each cloud, $N_{\rm H}=N_{\rm H,a}+N_{\rm H,b}$$\gamma_{\rm a}$ and $\gamma_{\rm b}$ are the magnetic inclination angles of the two clouds, and $\epsilon$ cancels out when calculating $p$.
@@
We divide Equations 12,  13, and 14 by $N_{\rm H,a}$ to be able to write them in terms of $\mathcal{N}_{c}$. Therefore,
$I/N_{\rm H,a}~{}=\epsilon\mathcal{N}_{c}-\frac{\alpha\epsilon}{2}\left[\cos^{2}{\gamma_{\rm a}}+(\mathcal{N}_{c}-1)\cos^{2}{\gamma_{\rm b}}-\frac{2}{3}\mathcal{N}_{c}\right],$ (16)
$Q/N_{\rm H,a}~{}=-\alpha\epsilon\left[\cos{2\psi_{a}}\cos^{2}{\gamma_{\rm a}}+(\mathcal{N}_{c}-1)\cos{2\psi_{b}}\cos^{2}{\gamma_{\rm b}}\right],$ (17)
$U/N_{\rm H,a}~{}=-\alpha\epsilon\left[\sin{2\psi_{a}}\cos^{2}{\gamma_{\rm a}}+(\mathcal{N}_{c}-1)\sin{2\psi_{b}}\cos^{2}{\gamma_{\rm b}}\right],$ (18)
where $N_{\rm H,a}$ cancels out when calculating $p$.
@@
We uniformly sample orientations in the range $[0,~{}\pi]$ for $\psi_{a}$ and $\psi_{b}$, values in the range $[0,~{}1]$ for $\cos{\gamma_{\rm a}}$ and $\cos{\gamma_{\rm b}}$, and a value in the range [1, 1.1] for $\mathcal{N}_{c}$ to calculate an instance of $p(\mathcal{N}_{c}\leq 1.1)$. We also sample different orientations and a value in the range [1.5, 2] for $\mathcal{N}_{c}$ to calculate an instance of $p(\mathcal{N}_{c}\geq 1.5)$. We then subtract $p(\mathcal{N}_{c}\leq 1.1)$ from $p(\mathcal{N}_{c}\geq 1.5)$ as in Section 4. We repeat this 10,000 times and average the results. We set $\alpha~{}=~{}0.22$, which corresponds to a maximum dust polarization fraction across the sky $p_{\rm max}~{}=~{}0.22$. This is the value Planck Collaboration et al. (2020a) observe for $p_{\rm max}$ at 353 GHz and 80 resolution. We get
$\langle p(\mathcal{N}_{c}~{}\geq~{}1.5)~{}-~{}p(\mathcal{N}_{c}~{}\leq~{}1.1)\rangle~{}=~{}-0.013.$ (19)
@@
This mean difference depends on the value for $\alpha$. For instance, if we set $\alpha~{}=~{}0.15$ as in Padoan et al. (2001) instead, we get a mean difference of -0.009. However, our estimate for the mean difference agrees with the result we measure in Section 4 for all reasonable values of $\alpha$. Thus, our empirical result is consistent with our theoretical estimate for the dust depolarization attributable to the line-of-sight dust complexity.
@@
Variations in the orientation of magnetic fields along the line of sight induce differences in the polarization angles of different dust components along the same sightline (Lee & Draine, 1985; Tassis & Pavlidou, 2015; King et al., 2018). When the emission of those components has different spectral energy distributions (SEDs), a frequency-dependent variation of the observed dust polarization angle along that sightline emerges, a phenomenon known as line-of-sight frequency decorrelation. This decorrelation complicates the translation of polarized dust emission maps from one frequency to another. Current analysis within the BICEP/Keck field does not demonstrate evidence of this phenomenon (BICEP/Keck Collaboration et al., 2021, 2023). However, a statistically significant detection of line-of-sight frequency decorrelation has been identified in larger sky areas across sightlines intersecting multiple dust clouds with varying magnetic field orientations (Pelgrims et al., 2021). Since polarized dust emission is the major foreground for CMB polarization measurements at high frequencies, it is important to characterize how the spatial complexity of the magnetic field in the dust might affect the frequency dependence of the foreground signal. In this paper, we have presented evidence that the 3D spatial complexity of the dust affects the dust polarization signal, even at a fixed frequency.
@@
The analysis in this paper highlights the importance of 3D dust mapping out to large distances. The dust distribution affects the dust polarization fraction, which has implications for the 3D magnetic field distribution. Since we expect higher complexity sightlines to have a lower dust polarization fraction on average, we would infer that the magnetic field is more uniform along a sightline if it has both a highly complex dust distribution and a large dust polarization fraction. This analysis was performed on the sightlines shown in Figure 6. Improvements to 3D dust modeling will allow us to look for this effect at the very high Galactic latitudes excluded here.